# Splitting of Girvin-MacDonald-Platzman density wave and the nature of chiral gravitons in fractional quantum Hall effect

### v1

@@ Results:
In all cases that we studied, it was found that the GMP mode, parton excitons, and CFEs constructed using direct projection at $L{=}1$ vanish identically. At higher $L$, we have studied the following ($\nu$, $N$) systems with direct projection: ($1/3$,6) for $L{=}2,3,4,5,6,7$; ($1/3$,8) for $L{=}2,3,4,5,6,7$; ($2/5$,6) for $L{=}2,3,4$; ($2/5$,8) for $L{=}2,3,4,5$; ($2/3$,6) for $L{=}2,3,4$; ($2/3$,8) for $L{=}2,3,4$; ($2/7$,6) for $L{=}2,3,4$; ($2/7$,8) for $L{=}2,3,4$; ($2/9$,6) for $L{=}2,3,4$; bosonic ($2/3$,6) for $L{=}2,3,4$; bosonic ($2/3$,8) for $L{=}2,3,4$. For all of these, we find ${\cal O}{=}1$ to machine precision, i.e. the GMP mode is exactly contained in the CFE space.
@@
SII GMP ansatz using the exact LLL Coulomb ground state

### v2

@@ Results:
In all cases that we studied, it was found that the GMP mode, parton excitons, and CFEs constructed using direct projection at $L{=}1$ vanish identically. At higher $L$, we have studied the following ($\nu$, $N$) systems with direct projection: ($1/3$,6) for $L{=}2,3,4,5,6,7$; ($1/3$,8) for $L{=}2,3,4,5,6,7$; ($2/5$,6) for $L{=}2,3,4$; ($2/5$,8) for $L{\leq}11$; ($2/3$,6) for $L{=}2,3,4$; ($2/3$,8) for $L{=}2,3,4$; ($2/7$,6) for $L{=}2,3,4$; ($2/7$,8) for $L{=}2,3,4$; ($2/9$,6) for $L{=}2,3,4$; ($2/9$,8) for $L{=}2$; bosonic ($2/3$,6) for $L{=}2,3,4$; bosonic ($2/3$,8) for $L{=}2,3,4$. For all of these, we find ${\cal O}{=}1$ in arbitrary precision calculations, i.e. the GMP mode is exactly contained in the CFE space.
@@
SIII GMP ansatz using the exact LLL Coulomb ground state
@@
SIV Effect of disorder broadening and resolution on the spectral function
The spectral function of the GMP mode (which forms the DSF) shown in Fig. 4 in the main text is estimated by replacing the delta functions appearing on the right-hand-side of Eq. (1) in the main text with normal distributions of width $\sigma$ as follows
$S(\mbox{\boldmath$q$},E)\approx\sum_{\alpha}|\langle\Psi_{\nu}^{\alpha}|\Psi^{\rm GMP}_{\mbox{\boldmath$q$}}\rangle|^{2}\frac{1}{\sqrt{2\pi\sigma}}\exp\left[\frac{(E_{\alpha}-E)^{2}}{2\sigma^{2}}\right].$ (S3)
Fig. S2 shows the effect of $\sigma$ on the estimated spectral function at filling $\nu{=}2/9$ of $N{=}6$ particles. The value of $\sigma$ may be thought of as arising from the disorder-broadening of the peaks or the resolution of the measurement. The spectral function shows two prominent peaks for low resolution, i.e. for values of $\sigma$ larger than the energy level spacing within the $L{=}2$ angular momentum sector.
Figure S2: Effect of bin-width on the spectral function of the GMP mode at $\nu{=}2/9$, $N{=}6$. As in Fig. 4 of the main text, the blue and red dots show spectral weights with the eigenstates obtained from diagonalization of Coulomb interaction in the full Hilbert space and CFE subspace respectively. The spectral function is estimated by replacing the delta functions in Eq. (1) of the main text with normal distributions of width $\sigma$. Different panels show the spectral functions calculated using different choices of the $\sigma$.
SV Dynamical structure factor at other angular momenta
Figure S3 shows the spectral weights and DSF for the GMP mode of the $\nu{=}2/5$ state for angular momentum sectors $L{=}2$ to $L{=}11$ for $N{=}8$ electrons. A single peak is visible at the lowest angular momentum $L{=}2$ as expected for a primary Jain sequence state, however, more peaks emerge at larger momenta. The DSF peaks appear at higher energies at larger momenta. The spectral functions forming obtained from calculations involving only the directed projected CFE space match the ones obtained by the calculations in the full Hilbert space till $L{\approx}9$ with the agreement being much better at lower angular momenta.
Figure S3: Spectral weights (dots) and spectral functions (continuous lines) for the GMP mode on the $\nu{=}2/5$ state obtained from full Hilbert space ED calculations (blue) and direct projected CFE subspace (red) shown for the angular momenta sectors $112{\leq}L{\leq}$ 11.